# Estimates of automorphic forms on $\mathrm{SU}(n,1)$

### v1

@@ 3.1. Estimates of the Bergman kernel
$\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}\ll\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+u_{2}^{2}+\sum_{j=2}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}+$
(51) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1/2}}\cdot\frac{\Gamma(k(n+1)-1)}{\Gamma(k(n+1)-1/2)},$
@@
$\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\ll$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-1}}+$
(52) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-3/2}}\cdot\frac{\Gamma(k(n+1)-3/2)}{\Gamma(k(n+1)-1)}.$
@@
$\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{2}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)}}+$
(53) $\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{2}\in\mathbb{Z}}\Bigg{(}\sum_{\alpha=0}^{2}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)-\alpha/2}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-(\alpha+1)/2}}\Bigg{)}.$
@@
$\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{2}\in\mathbb{Z}}\Bigg{(}\sum_{\alpha=0}^{2}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}2x_{1}\big{)}^{k(n+1)-\alpha/2}}{\big{(}2x_{1}+\sum_{j=2}^{n}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-(\alpha+1)/2}}\Bigg{)}=$
(56) $\hskip 233.3125ptO_{X_{\Gamma}}\big{(}\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sqrt{x}\big{)}=O_{X_{\Gamma}}\big{(}k^{n+1/2}\big{)}.$
@@ 3.2. Estimates of the Bergman metric
With notation as above, for $k\gg 1$ and $z\in\overline{X}_{\Gamma}$, we have the following lower bound
(60) $\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}\gg_{\Gamma}k^{n},$
where the implied constant depends only on $X_{\Gamma}$.
@@
As in the proof of Theorem 3.4, for a fixed $\epsilon_{0}>0$, which is independent of $k$, let $\mathcal{A}_{\epsilon_{0}}:=\mathcal{F}_{\Gamma_{3}}\backslash U_{\epsilon_{0}}(\infty)$ be a fixed compact subset. For $k\gg 1$, and $z\in\mathcal{A}_{\epsilon_{0}}$, adapting arguments from the proof of Proposition 3.2, we have the following estimate
$\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(\gamma_{31}z)\big{|}_{\mathrm{pet}}\gg_{\Gamma}\sum_{\gamma\in\Gamma_{3}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)}\gg_{\Gamma}$
(61) $\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}-\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\epsilon_{0}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\epsilon_{0}}/4)}-\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\epsilon_{0}}/8)}{\sinh^{2n}(r_{\epsilon_{0}}/4)\cosh^{k(n+1)}(3r_{\epsilon_{0}}/8)}\gg_{\Gamma}k^{n},$
$\mathrm{where}\,\,r_{\epsilon_{0}}:=\inf\big{\{}\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)\big{|}\,z\in\mathcal{A}_{\epsilon_{0}},\,\,\gamma\in\Gamma_{3}\backslash\{\mathrm{Id}\}\big{\}}.$
For $z\in U_{\epsilon_{0}}(\infty)$, from the proof of Theorem 3.4, we have
$\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(\gamma_{31}z)\big{|}_{\mathrm{pet}}\geq\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}+\Bigg{|}\sum_{\gamma\in\Gamma_{3,\infty}\backslash\{\mathrm{Id}\}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\langle z,z\rangle_{H_{3}}^{k(n+1)}}{\langle z,\gamma z\rangle_{H_{3}}^{k(n+1)}}\Bigg{|}-$
@@
(62) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(\operatorname{{\it{r_{X_{\mathrm{\Gamma}}}}}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(\operatorname{{\it{r_{X_{\mathrm{\Gamma}}}}}}/4)}-\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5\operatorname{{\it{r_{X_{\mathrm{\Gamma}}}}}}/8)}{\sinh^{2n}(\operatorname{{\it{r_{X_{\mathrm{\Gamma}}}}}}/4)\cosh^{k(n+1)}(3\operatorname{{\it{r_{X_{\mathrm{\Gamma}}}}}}/8)}.$
For $z\in U_{\epsilon_{0}}(\infty)\backslash\partial U_{4\pi/k(n+1)}(\infty)$, and from the description of elements of $\Gamma_{3,\infty}$, which is as described in equation (12), we infer that each term of the series
@@
$\Bigg{|}\sum_{\gamma\in\Gamma_{3,\infty}\backslash\{\mathrm{Id}\}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\langle z,z\rangle_{H_{3}}^{k(n+1)}}{\langle z,\gamma z\rangle_{H_{3}}^{k(n+1)}}\Bigg{|}$
approaches zero, as $k$ approaches $\infty$. So we can conclude that
$\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(\gamma_{31}z)\big{|}_{\mathrm{pet}}\gg_{\Gamma}\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\gg_{\Gamma}k^{n}.$
For $z\in\partial U_{4\pi/k(n+1)}(\infty)$, from the proof of Theorem 3.4, we deduce that
@@
$\Bigg{|}\sum_{\gamma\in\Gamma_{3,\infty}\backslash\{\mathrm{Id}\}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\langle z,z\rangle_{H_{3}}^{k(n+1)}}{\langle z,\gamma z\rangle_{H_{3}}^{k(n+1)}}\Bigg{|}\gg_{\Gamma}k^{n},$
which completes the proof of the proposition. ∎
@@
With notation as above, for $k\gg 1$, and $1\leq i,j\leq n$, and a fixed $a\in\mathbb{R}$ with $-2n\leq a\leq 2n$, we have the following estimate
@@
Let $\{f_{1},\ldots,f_{d_{k}}\}$ denote an orthonormal basis of $\operatorname{{\it{S_{k(n+1)}}}(\Gamma)}$. For any $z\in\mathcal{F}_{\Gamma}$, and $1\leq i,j\leq n$, and a fixed $a\in\mathbb{R}$ with $-2n\leq a\leq 2n$, from the definition of the Bergman kernel, and the Cauchy-Schwartz inequality, we have
@@
With notation as above, for $k\gg 1$, and $1\leq i\leq n$, and a fixed $a\in\mathbb{R}$ with $-2n\leq a\leq 2n$, , we have the following estimate
@@
Let $\{f_{1},\ldots,f_{d_{k}}\}$ denote an orthonormal basis of $\operatorname{{\it{S_{k(n+1)}}}(\Gamma)}$. For any $z\in\mathcal{F}_{\Gamma}$, and $1\leq i,j\leq n$, and a fixed $a\in\mathbb{R}$ with $-2n\leq a\leq 2n$, from the definition of the Bergman kernel, and Cauchy-Schwartz inequality, we have
@@
With notation as above, for $k\gg 1$, and $1\leq i\leq n$, and a fixed $a\in\mathbb{R}$ with $-2n\leq a\leq 2n$, we have the following estimate
@@
(73) $\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\operatorname{\mu_{\mathrm{Ber},{\it{k}}}^{\mathrm{vol}}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{4n-1}\big{)},$
@@
From Lemma 3.5, and from the formula for the hyperbolic volume form, which as described in equation (6), we have
@@
[…]
For any $a\in\mathbb{R}$ with $-n\leq a\leq n$, combining Propositions 3.6 and 3.7, we arrive at the following estimate
@@
(76) $\frac{\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}}\cdot\bigg{|}\frac{\partial^{2}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}\partial\overline{z}_{\beta}}\bigg{|}O_{X_{\Gamma}}\big{(}k^{a+9/2}\big{)}.$

### v2

@@ 3.1. Estimates of the Bergman kernel
$\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}\ll\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)}}+$
(52) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-1/2}}\cdot\frac{\Gamma(k(n+1)-1)}{\Gamma(k(n+1)-1/2)},$
@@
$\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\ll$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-1}}+$
(53) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-3/2}}\cdot\frac{\Gamma(k(n+1)-3/2)}{\Gamma(k(n+1)-1)}.$
@@
$\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{2}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)}}+$
(54) $\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{2}\in\mathbb{Z}}\Bigg{(}\sum_{\alpha=0}^{2}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)-\alpha/2}}{\big{(}k(n+1)/4\pi+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-(\alpha+1)/2}}\Bigg{)}.$
@@
$\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{2}\in\mathbb{Z}}\Bigg{(}\sum_{\alpha=0}^{2}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}2x_{1}\big{)}^{k(n+1)-\alpha/2}}{\big{(}2x_{1}+\sum_{j=2}^{n-1}(u_{j}^{2}+v_{j}^{2})+u_{1}^{2}\big{)}^{k(n+1)-(\alpha+1)/2}}\Bigg{)}=$
(57) $\hskip 233.3125ptO_{X_{\Gamma}}\big{(}\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sqrt{x_{1}}\big{)}=O_{X_{\Gamma}}\big{(}k^{n+1/2}\big{)}.$
@@ 3.2. Estimates of the Bergman metric
With notation as above, for $k\gg 1$ and $z=(z_{1},\ldots,z_{n})^{t}\in U_{2\pi/(k(n+1))}(\infty)$, we have the following lower bound
(61) $\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}\gg_{X_{\Gamma}}\frac{k^{n}}{\sqrt{-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}}};$
when $z=(z_{1},\ldots,z_{n})^{t}\in\mathcal{F}_{\Gamma_{3}}\backslash U_{2\pi/(k(n+1))}(\infty)$, we have the following lower bound
(62) $\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}\gg_{X_{\Gamma}}\bigg{(}\frac{-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}}{k}\bigg{)}^{(n-1)/2}k^{n-1/2},$
and the implied constants in both estimates depends only on $X_{\Gamma}$.
@@
As in the proof of Theorem 3.4, for any $z\in\mathcal{F}_{\Gamma_{3}}$, we have the following estimate
@@
$\sum_{\gamma\in\Gamma_{3}\backslash\Gamma_{3,\infty}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)}\ll_{X_{\Gamma}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{X_{\Gamma}}/8)}{\sinh^{2n}(r_{X_{\Gamma}}/4)\cosh^{k(n+1)}(3r_{X_{\Gamma}}/8)}+$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{X_{\Gamma}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{X_{\Gamma}}/4)\cosh^{k(n+1)-2n-2}(3r_{X_{\Gamma}}/8)}\ll_{X_{\Gamma}}\frac{k^{n}}{\cosh^{k(n+1)}(3r_{X_{\Gamma}}/8)},$
(63) $\mathrm{where}\,\,r_{X_{\Gamma}}:=\inf\big{\{}\mathrm{d}_{\mathrm{hyp}}(z,\gamma z)\big{|}\,z\in\mathcal{F}_{\Gamma_{3}},\,\gamma\in\Gamma_{3}\backslash\Gamma_{3,\infty}\big{\}}.$
With notation as in Theorem 3.4, and using equation (3.4), we have
@@
(64) $\sum_{\gamma\in\Gamma_{3,\infty}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\langle\tilde{z},\tilde{z}\rangle_{H_{3}}^{k(n+1)}}{\langle\tilde{z},\tilde{\gamma z}\rangle_{H_{3}}^{k(n+1)}}=\sum_{(\tau,L)\in L}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}\big{)}^{k(n+1)}}{\big{(}-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}+\frac{|\tau|^{2}}{2}+i\frac{t}{2}\big{)}^{k(n+1)}}.$
Put
$\alpha(z):=-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}.$
From Remark 2.1, and using Poisson summation formula, we have
$\sum_{(\tau,L)\in L}\frac{\big{(}-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}\big{)}^{k(n+1)}}{\big{(}-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}+\frac{|\tau|^{2}}{2}+i\frac{t}{2}\big{)}^{k(n+1)}}\gg_{X_{\Gamma}}\sum_{\tau\in\mathbb{Z}^{2n-2}}\sum_{t\in\mathbb{Z}}\frac{\alpha(z)^{k(n+1)}}{\big{(}\alpha(z)+|\tau|^{2}+it\big{)}^{k(n+1)}}=$
(65) $\sum_{\tau\in\mathbb{Z}^{2n-2}}\sum_{m\in\mathbb{Z}_{\geq 1}}\frac{2\pi\alpha(z)^{k(n+1)}e^{-2\pi m(|\tau|^{2}+\alpha(z))}(2\pi m)^{k(n+1)}}{(k(n+1)-1)!}.$
Using integral test for the lattice $\tau\in\mathbb{Z}^{2n-2}$, we derive
$\sum_{\tau\in\mathbb{Z}^{2n-2}}\sum_{m\in\mathbb{Z}_{\geq 1}}\frac{2\pi\alpha(z)^{k(n+1)}e^{-2\pi m(|\tau|^{2}+\alpha(z))}(2\pi m)^{k(n+1)}}{(k(n+1)-1)!}\gg_{X_{\Gamma}}$
$\sum_{m\in\mathbb{Z}_{\geq 1}}\bigg{(}\int_{0}^{\infty}e^{-2\pi mx^{2}}dx\bigg{)}^{2n-2}\cdot\frac{2\pi\alpha(z)^{k(n+1)}e^{-2\pi m\alpha(z)}(2\pi m)^{k(n+1)}}{(k(n+1)-1)!}\gg_{X_{\Gamma}}$
(66) $\sum_{m\in\mathbb{Z}_{\geq 1}}\frac{(2\pi\alpha(z))^{k(n+1)}m^{k(n+1)-(n+1)/2}e^{-2\pi m\alpha(z)}}{(k(n+1)-1)!}.$
Using integral test for the summation on $m\in\mathbb{Z}_{\geq}1$, we now compute
$\sum_{m\in\mathbb{Z}_{\geq 1}}\frac{(2\pi\alpha(z))^{k(n+1)}m^{k(n+1)-(n+1)/2}e^{-2\pi m\alpha(z)}}{(k(n+1)-1)!}\geq$
$\frac{(2\pi\alpha(z))^{(n+1)/2}}{(k(n+1)-1)!}\int_{1}^{\infty}(2\pi m\alpha(z))^{k(n+1)-(n+1)/2}e^{-2\pi\alpha(z)}dm\geq$
(67) $\frac{(2\pi\alpha(z))^{(n-1)/2}}{(k(n+1)-1)!}\int_{2\pi\alpha(z)}^{\infty}\theta(z)^{k(n+1)-(n+1)/2}e^{-\theta}d\theta.$
For $z=(z_{1},\ldots,z_{n})^{t}\in U_{2\pi/(k(n+1))}(\infty)$, since $2\pi\alpha(z)>k(n+1)$, combining estimates (64)–(3.6) with Proposition 3.2, using Sterling’s approximation, we have the following lower bound
@@
(68) $\sum_{\gamma\in\Gamma_{3,\infty}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\langle\tilde{z},\tilde{z}\rangle_{H_{3}}^{k(n+1)}}{\langle\tilde{z},\tilde{\gamma z}\rangle_{H_{3}}^{k(n+1)}}\gg_{X_{\Gamma}}\frac{k^{n}}{\sqrt{-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}}}.$
Similarly, for $z=(z_{1},\ldots,z_{n})^{t}\in\mathcal{F}_{\Gamma_{3}}\backslash U_{2\pi/(k(n+1))}(\infty)$, since $k(n+1)<2\pi\alpha(z)$, combining estimates (64)–(3.6) with Proposition 3.2, using Sterling’s approximation, we have the following lower bound
(69) $\sum_{\gamma\in\Gamma_{3,\infty}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\langle\tilde{z},\tilde{z}\rangle_{H_{3}}^{k(n+1)}}{\langle\tilde{z},\tilde{\gamma z}\rangle_{H_{3}}^{k(n+1)}}\gg_{X_{\Gamma}}\bigg{(}\frac{-2\mathrm{Re}(z_{1})-\sum_{i=2}^{n}|z_{i}|^{2}}{k}\bigg{)}^{(n-1)/2}k^{n-1/2},$
For $k\gg 1$, combining estimate (3.6) with estimates (68) and (69), completes the proof of the proposition. ∎
@@
With notation as above, for $k\gg 1$, and $1\leq i,j\leq n$, and a fixed $a\in\mathbb{R}$ with $2n\leq a\leq 2n$, we have the following estimate
@@
Let $\{f_{1},\ldots,f_{d_{k}}\}$ denote an orthonormal basis of $\operatorname{{\it{S_{k(n+1)}}}(\Gamma)}$. For any $z\in\mathcal{F}_{\Gamma}$, and $1\leq i,j\leq n$, and a fixed $a\in\mathbb{R}$ with $2n\leq a\leq 2n$, from the definition of the Bergman kernel, and the Cauchy-Schwartz inequality, we have
@@
With notation as above, for $k\gg 1$, and $1\leq i\leq n$, and a fixed $a\in\mathbb{R}$ with $2n\leq a\leq 2n$, , we have the following estimate
@@
Let $\{f_{1},\ldots,f_{d_{k}}\}$ denote an orthonormal basis of $\operatorname{{\it{S_{k(n+1)}}}(\Gamma)}$. For any $z\in\mathcal{F}_{\Gamma}$, and $1\leq i,j\leq n$, and a fixed $a\in\mathbb{R}$ with $2n\leq a\leq 2n$, from the definition of the Bergman kernel, and Cauchy-Schwartz inequality, we have
@@
With notation as above, for $k\gg 1$, and $1\leq i\leq n$, and a fixed $a\in\mathbb{R}$ with $2n\leq a\leq 2n$, we have the following estimate
@@
(80) $\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\operatorname{\mu_{\mathrm{Ber},{\it{k}}}^{\mathrm{vol}}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{2(n-1)(n+2)+n+3}\big{)},$
@@
For any $z=(z_{1},\ldots,z_{n})^{t}\in\mathcal{F}_{\Gamma}$, from Lemma 3.5, and from the formula for the hyperbolic volume form, which as described in equation (6), we have
@@
and $\mathcal{I}\subset\big{\{}(i,j)|\,1\leq i,j\leq n\}$, and $|\mathcal{I}|$, the cardinality of the set $\mathcal{I}$ is equal to nln-l.
For any $a\in\mathbb{R}$ with $n\leq a\leq n$, from Propositions 3.6 (we consider estimate (62), as the contribution of estimate (62) dominates the contribution of estimate (61)), and using the definition of map $\gamma_{13}$, for any $z\in\mathcal{F}_{\Gamma}$ we arrive at the following estimate
@@
$\frac{\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(\gamma_{13}z)\big{|}_{\mathrm{pet}}}\cdot\bigg{|}\frac{\partial^{2}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}\partial\overline{z}_{\beta}}\bigg{|}\ll_{X_{\Gamma}}\frac{k^{-n/2}\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\alpha(\gamma_{13}z)\big{|}^{(n-1)/2}}\cdot\bigg{|}\frac{\partial^{2}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}\partial\overline{z}_{\beta}}\bigg{|},$
where
$\alpha(\gamma_{13}z)=\frac{2(1-|z|^{2})}{|z_{1}-1|^{2}}.$
Combining the above estimate with Proposition 3.7, we arrive at the following esimate
$\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}\frac{\partial^{2}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}\partial\overline{z}_{\beta}}\bigg{|}\ll_{X_{\Gamma}}$
(83) $k^{-n/2}\big{(}1-|z|^{2}\big{)}^{k(n+1)-(n-1)/2-a}\bigg{|}\frac{\partial^{2}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}\partial\overline{z}_{\beta}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{n+a+4}\big{)}.$